# Qualitative Study for LLM-assisted Design Study Process: Strategies, Challenges, and Roles

### v2

@@ 4 Qualitative Interview Results
In this section, we present the qualitative feedback collected from 30 participants, focusing on their experiences with leveraging LLMs to assist design studies in individual stages of the Sedlmair et al. framework [34]. We explore the strategies they employed, the challenges they encountered, and any useful practices they adopted in the study process. We label findings with the associated LLM roles. Our analysis aims to uncover diverse insights that emerge across participants with varying levels of visualization expertise and LLM usage frequency.
@@ 1 Precondition: Learn
“In summary, LLMs were found to be highly effective in aiding novice researchers to quickly grasp domain fundamentals, while more advanced users leverage them to stay updated with recent developments, together with a “validation-loop” between visualization researchers, domain experts, and LLMs. However, challenges such as inaccuracies in technical details and confirmation bias need users to take practices like consulting domain experts for reliability.”
@@
The “Learn” stage highlights how researchers use LLMs to quickly obtain foundational knowledge in unfamiliar domains or stay updated with recent developments in their own fields. Most novice visualization developers found LLMs particularly helpful for quickly grasping fundamental concepts in a specific domain connector. For example, P6 noted that LLMs allowed them to “quickly grasp the basics of a domain that would otherwise require extensive manual searching.” However, proficient and advanced users use LLMs to acquire the latest developments in their fields.
@@
Despite these benefits, participants faced challenges, particularly novices, who often treated LLMs as search engines without the expertise to discern inaccuracies in the responses. P5 mentioned that “LLMs occasionally provide incorrect answers, but as a beginner, it is difficult to tell if there are actually any incorrect contents.” To overcome this uncertainty, novices are often forced to double-check information with people with more expertise at a significant cost of time. Another common challenge was the accuracy of detailed explanations provided by LLMs. While their high-level summaries were generally correct, technical details were sometimes erroneous, as noted by P23, who gave an example where the explanation of an object detection algorithm was conceptually accurate, but the technical details about the model’s trainable layer were incorrect. To address these issues, participants adopted various practices, including consulting domain experts, cross-referencing LLM outputs with literature, and generating tailored learning materials using LLMs. For instance, P9 used LLMs to create coding tutorials and background explanations for self-training in new domains. But he noted the limitations that LLMs could not effectively use many emerging but immature tools (e.g., the new package ‘turf.js’ in the field of geospatial visualization), even when provided with its tutorial.
@@
Another interesting finding is from advanced users. They mentioned a “validation loop”, involving a dynamic process where visualization researchers, domain experts, and LLMs cross-verify with each other connector assistant. For example, visualization researchers used LLMs to validate domain knowledge provided by domain practitioners, while domain experts reviewed LLM-generated designing methods with visualization researchers to ensure alignment with real practice. Meanwhile, visualization users also validate the LLM-generated domain knowledge with domain practitioners as well. P21 described this triangulation process as effective for ensuring reliability but sometimes introducing pitfalls, because LLMs often aligned with the phrasing embedded in prompts, a phenomenon referred to as “confirmation bias.” One example is that if the domain experts ask LLM “Please help me to decide if the stacked bar charts can fit my data or not,” and LLM will most likely provide you with the answers about why stacked bar charts can fit your data instead of critical answers. To mitigate this, P21 recommended allowing LLMs to generate independent evidence and conclusions without bias.
@@
Furthermore, participants also highlighted unique strategies for leveraging LLMs during this “Learn” stage. For example, P24 used LLMs to fill gaps in domain-specific terminology connector, enabling them to bridge knowledge gaps more efficiently. However, P21 also warned against relying on LLMs for summarizing design requirements, as these summaries often failed to align with their domain expertise. Similarly, P9 employed LLMs to generate tailored courses and answers for self-training for a new domain and also the technical stack connector programmer.
@@ 2 Precondition: Winnow
The “Winnow” stage is about narrowing down domain problems and collaborators. Novice students often used LLMs to evaluate whether visualization techniques could effectively solve a given problem. Specifically, they prompted LLMs to provide the most relevant non-visualization papers to check if there were existing automatic approaches or algorithms that could address the problem assistant. Additionally, they explored whether visualization methods, even from different areas, could be generalized and applied to their contexts assistant. However, a major challenge arose due to LLM hallucinations, as popular models frequently recommended papers that did not exist. To overcome this, participants turned to LLM-powered literature search platforms, such as Scite, to conduct more reliable literature reviews and verify the authenticity of suggested resources. PhD students noted that prior to the availability of LLMs, they primarily relied on personal connections or pre-existing knowledge of researchers in the domain for collaborator recruitment. With LLMs, however, they could directly request a list of relevant researchers working in the same field assistant. Subsequently, some advanced researchers utilized LLMs to generate recruitment emails for inviting collaborators from other domains assistant.
While this approach saved time, it introduced specific challenges. First, the emails were often too generic, lacking personalized details such as how their research aligned with the project. To address this, participants attached the recipient’s CV or homepage link to the LLM prompt, enabling the model to generate more tailored emails. Second, privacy concerns arose when describing research projects to LLMs, as the model could extract sensitive keywords or overly detailed descriptions, risking unintentional idea leakage. To mitigate this, participants often manually verify generated content before sharing.
@@ 3 Precondition: Cast
The “Cast” stage focuses on forming the right team of collaborators for a design study. According to our interview, advanced beginners in visualization used LLMs to assign roles within a project by providing the LLMs with collaborators’ CVs or brief descriptions of their expertise assistant. While this approach facilitates the initial role assignment, participants noted challenges in achieving enough granularity. For example, it was difficult for LLMs to distinguish whether a collaborator with a visualization background would be better suited as a designer or a programmer. Furthermore, LLMs lacked the ability to assess personal traits, such as leadership or communication skills, making it challenging to identify appropriate candidates for roles like team managers or connectors, who are the key members responsible for coordinating all members. To address these limitations, participants often relied on the Principal Investigator (PI) or a designated “gatekeeper” to make the final decisions regarding role assignments.
Fresh PhD students faced additional difficulties in distinguishing between front-line analysts and translators within their teams connector. Thus, they usually used LLMs to analyze collaborators’ published papers and professional backgrounds, allowing the model to suggest suitable candidates for these roles. Beyond role assignment, LLMs can also be directly regarded as an “agent” to facilitate collaboration among researchers from different domains connector. For example, LLMs helped by providing task abstractions and translating domain-specific terminology into language that visualization researchers could understand.
@@ 4 Core: Discover
The “Discover” stage highlights how researchers use LLMs to uncover knowledge, identify related work, and understand domain-specific requirements. Novice researchers often used LLMs to explore the literature review by asking LLMs to suggest papers or summarize the latest research trends assistant connector, which was particularly helpful for those unfamiliar with the domain. However, this strategy often introduced hallucinated references to novices, where LLMs usually provided nonexistent papers or incorrect publication details. According to P21, who is a professor and lecturer in visualization, LLMs often fail to recognize the graphical visual design elements, but the students cannot totally identify this incorrectness. Proficient PhD students and professors used LLMs to summarize the takeaways from the academic paper assistant, e.g., key contributions, methodologies, and findings from visualization papers. However, participants commented that LLMs often misinterpreted the contributions of visualization papers, particularly related to the visual design elements. For instance, the models often failed to accurately understand visual encodings or the rationale behind specific visual channels. This limitation required significant additional manual effort to validate outputs.
Beyond literature discovery, participants also explored interesting ways to use LLMs to understand domain-specific requirements connector. Instead of relying solely on interviews with domain experts, some researchers used LLMs to analyze routine tasks performed by professionals in their fields. For example, P12 described using Copilot embedded in an IDE to document how bioinformatics researchers analyze tumor tissue slices, which captures the practical workflows that might otherwise be overlooked in traditional interviews. This approach can significantly reduce information loss and provide a more precise understanding of domain requirements.
Participants also expressed that they often use LLMs to double-check the collected design requirements from domain experts connector. However, LLMs’ answers are usually too generic to help with the requirement validation. This limitation was particularly evident when researchers leveraged LLMs to generate design requirements or propose novel solutions. As a result, researchers emphasized the importance of combining LLM outputs with their own expertise, using the models as tools to augment, rather than replace, their discovery process. P14 also mentioned the usage of the generation of the questionnaire for the pilot study or formative study assistant. Also, P14 used LLMs to generate the possible answers from the provided questionnaire simulator, which is used to compare with the real answers from the participants to identify some unique and insightful findings in this domain.
@@ 5 Core: Design
The “Design” stage focuses on leveraging LLMs to assist visualization researchers in enhancing the proposed visual design, global layout, user interaction, and prototype usability. Three professors highlighted the use of LLMs as “translators” during the design process to bridge the gap between visualization designers and domain experts. In this case, LLMs were used to facilitate communication by summarizing iterative meetings, generating meeting notes, and translating domain-specific knowledge and terminology into practical insights for visualization designers connector. Some PhD students reported using LLMs to decompose domain requirements, obtained during the “Discover” stage, into sub-tasks and brainstorm design ideas based on these sub-tasks assistant. For example, they would ask LLMs to search approaches used in prior visualization studies, such as those published in IEEE TVCG papers, for similar design requirements.
@@
For system-wide design tasks, such as determining the overall layout, interaction logic, and workflow between views, LLMs were used to provide high-level suggestions based on the screenshot together with the code simulator. However, participants noted that LLMs often struggled with these tasks because LLMs lacked the ability to maintain a comprehensive understanding of the entire system. This limitation required researchers to manually refine and integrate the LLM’s outputs into a cohesive global design.
LLMs were also employed as sources of design inspiration assistant. Some researchers provided a domain and asked the LLM to suggest straightforward visualization designs to inspire their work. While this approach often generated initial ideas, participants noted that the suggestions were often limited to those commonly-used approaches. To address this challenge, researchers came up with practices such as incorporating evaluation metrics (e.g., significance or novelty) to assess the quality of the proposed designs. Moreover, PhD designers used LLMs to search commercial tools online for existing solutions as inspiration for their designs.
@@ 6 Core: Implement
The “Implement” stage in the design study methodology explores how LLMs can be utilized to support various phases of visualization system development. According to our interview, this stage is the common scenario where visualization developers rely on LLMs. First of all, for data processing tasks, novices and PhD students can use LLMs to generate code for data cleaning or other pre-processing tasks programmer. Also, after the system prototype becomes mature, LLMs can be employed to generate large-scale datasets to test the system’s scalability programmer, which can support researchers to initially evaluate if the prototype can seamlessly fit the domain’s real needs.
In terms of technical stack setup, according to the findings from the professors, LLMs can help beginners to generate basic but functional architectures programmer, such as a simple React.JS and Flask framework that supports fundamental tasks like data requesting. This approach allows developers to focus on design implementation while avoiding the time-consuming process of learning new frameworks.
For implementing the design, LLMs can assist in generating code for visual design implementation programmer. While LLMs work well for basic chart designs when provided with a sketch or bitmap, they often fail to recognize hierarchical structures. For example, in the case of a stacked bar chart, LLMs may generate separate bars instead of a cohesive stacked structure. As a workaround, novices may turn to user-friendly tools like Tableau, but it introduced new issues of flexibility and interactivity, limiting the visualization to only simple analysis tasks. Another significant challenge arises when the generated JavaScript code is only suitable for static HTML pages but fails to integrate into component-based frameworks like React.JS. This incompatibility comes from several issues: complex interaction logic makes the code difficult to adapt, conflicts such as class name duplication can cause system bugs, the generated design may not align with the overall size or color scheme of the system, and SVG elements may not be correctly grouped. The best practices to address these challenges include providing detailed context about all components to hint LLMs and reduce integration issues.
Several challenges also exist in the process of fine-tuning prototype systems. For example, it is difficult for LLMs to modify specific annotated parts of a design, as they struggle to recognize which part of the design is being referenced. Also, if a user requests modifying the “top left” circle, LLMs lack spatial awareness and cannot identify the target. Moreover, some descriptions, such as making a color scheme “look cool” or “minimalist but not simplistic”, are often misunderstood by LLMs. A possible practice to address these issues is to write detailed comments in the code that explicitly specify the modifications required.
@@ 7 Core: Deploy
In the Deploy stage, LLMs are used to evaluate the system’s usability, explore potential usage scenarios, and test its performance under real-world constraints. Advanced students have reported using LLMs to simulate user behavior, allowing the model to “pretend” to be a user simulator. This approach helps uncover hidden and potential use cases for the system, as well as identify alternative workflows that differ from conventional ones proposed by human experts. However, two significant challenges arise: first, the generic background knowledge of LLMs can lead to superficial stories that lack depth; second, hallucinations by LLMs may result in incorrect stories based on non-existent system features.
@@
Moreover, proficient students have mentioned some useful ways to support user studies and feedback analysis by LLMs. Specifically, developers can transcribe user study recordings into text and use LLMs to summarize key takeaways and feedback assistant. Similarly, LLMs can assist in generating questionnaires and designing tasks based on the provided workflow and requirements. However, challenges exists in these practices. For example, relying on LLMs to summarize feedback may result in conclusions that do not align with the developer’s domain expertise. Additionally, when generating questionnaires based on historical papers, LLMs may over-rely on the provided materials, like previous papers, leading to repetitive questions.

### v4

@@ 4 Qualitative Interview Results
In this section, we present the qualitative feedback collected from 30 participants, focusing on their experiences with leveraging LLMs to assist design studies in individual stages of the Sedlmair et al. framework [34]. We explore the strategies they employed, the challenges they encountered, and any useful practices they adopted accordingly. To explicitly indicate clearly identify the roles introduced above, we assign labels to findings by placing the corresponding LLM role name adjacent to each strategy. Our analysis aims to uncover diverse practical insights in detail that emerge across participants with varying levels of visualization expertise and LLM usage frequency.
@@ 1 Precondition: Learn
“In summary, LLMs were found to be highly effective in aiding novice researchers to quickly grasp domain fundamentals, while more advanced users leverage them to stay updated with recent developments, together with a “validation loop” between visualization researchers, domain experts, and LLMs. However, challenges such as inaccuracies in technical details and confirmation bias need users to take practices like consulting domain experts for reliability.”
@@
The “Learn” stage indicates the need to acquire a solid understanding of visualization literature and techniques to inform the design study effectively. Most novice visualization developers found LLMs particularly helpful for quickly grasping fundamental concepts in a specific domain connector that would otherwise require extensive manual searching. However, proficient and advanced users use LLMs to acquire the latest developments in their fields in a short period of time.
@@
Despite these benefits, participants actually faced challenges, particularly novices, who often treated LLMs as search engines without the expertise to discern inaccuracies in the responses. For example, P5 mentioned that “LLMs occasionally provide incorrect answers, but as a beginner, it is almost impossible to tell if there are actually any incorrect contents.” To overcome this uncertainty, novices are often forced to double-check information with people with more expertise with a significant cost of time. Another common challenge was the accuracy of detailed explanations provided by LLMs. While their high-level summaries were generally correct, technical details were sometimes erroneous, as noted by P23, who gave an example where the explanation of an object detection algorithm was conceptually accurate, but the technical details regarding the model’s trainable layer were incorrect. To address these issues, participants adopted various practices, including consulting domain experts and cross-referencing LLM outputs with literature connector.
@@
Another interesting finding is from advanced users. They noted a “validation loop”, involving a dynamic process where visualization researchers, domain experts, and LLMs cross-verify with each other connector assistant. For example, visualization researchers used LLMs to validate domain knowledge provided by domain practitioners, while domain experts reviewed LLM-generated designing methods with visualization researchers to ensure alignment with real practice. Meanwhile, visualization users also validate the LLM-generated domain knowledge with domain practitioners. P21 described this triangulation process as effective for ensuring reliability but sometimes introducing pitfalls, because LLMs often aligned with the phrasing embedded in prompts. One example is that if the domain experts ask LLM “Please help me to decide if the stacked bar charts can fit my data or not,” and LLM will most likely provide you with the answers about why stacked bar charts “can” fit your data instead of critical answers. To mitigate this, P21 recommended allowing LLMs to generate independent evidence and conclusions without bias.
@@
Furthermore, participants also highlighted some unique strategies. P24 used LLMs to fill gaps in domain-specific terminology connector, enabling them to bridge knowledge gaps more efficiently. However, P21 also warned against relying on LLMs for summarizing design requirements, as these summaries often failed to align with their domain expertise. Similarly, P9 used LLMs to create coding tutorials and background explanations for self-training in new domains connector programmer. But he noted the limitations that LLMs could not effectively use many emerging but immature tools, even when provided with its detailed tutorial.
@@ 2 Precondition: Winnow
The “Winnow” stage is about selecting promising collaborations with domain experts by filtering out unsuitable partnerships early in the process. Specifically, novice students often used LLMs to evaluate whether visualization techniques could actually solve a given problem assistant. They prompted LLMs to provide the most relevant non-visualization papers to check if there were existing automatic approaches or algorithms that could address the problem. Additionally, they explored whether visualization methods, even from different areas, could be generalized and further applied into their contexts. However, a major challenge arose due to LLM hallucinations, as popular models frequently recommended papers that did not exist. To overcome this, participants turned to commercial LLM-powered literature searching tools to conduct more reliable literature reviews and verify the authenticity of suggested resources. PhD students noted that prior to the availability of LLMs, they primarily relied on personal connections or pre-existing knowledge of researchers in the domain for collaborator recruitment. With LLMs, however, they could directly request a list of relevant researchers working in the same field assistant.
Subsequently, some advanced researchers utilized LLMs to generate recruitment emails for inviting collaborators from other domains assistant. While this approach saved time, it introduced specific challenges. First, the emails were often too generic, lacking personalized details such as how their research aligned with the project. To address this, participants attached the recipient’s CV or homepage link to the LLM prompt, enabling the model to generate more tailored emails. Second, privacy concerns arose when describing research projects to LLMs, as the model could extract sensitive keywords or overly detailed descriptions, risking unintentional idea leakage. To mitigate this, participants often have to manually verify generated content before sharing.
@@ 3 Precondition: Cast
The “Cast” stage focuses on clarifying the roles of collaborators, ensuring that key participants, such as gatekeepers, are identified before starting the project. According to our interview, advanced beginners in visualization used LLMs to assign roles within a project by providing the LLMs with collaborators’ CVs or brief descriptions of their expertise assistant. However, participants noted challenges in achieving enough granularity. For example, it was difficult for LLMs to distinguish whether a collaborator with a visualization background would be better suited as a designer or a programmer. Furthermore, LLMs lacked the ability to assess personal traits, such as leadership or communication skills, making it challenging to identify appropriate candidates for roles like team managers or connectors, who are the key members responsible for coordinating all members. To address these limitations, participants often relied on the Principal Investigator (PI) or a designated “gatekeeper” to make the final decisions regarding role assignments.
Fresh PhD students faced additional difficulties in distinguishing between front-line analysts and translators within their teams. Thus, they usually used LLMs to analyze collaborators’ published papers and professional backgrounds, allowing the model to suggest suitable candidates for these roles connector. Beyond role assignment, LLMs can also be directly regarded as an “agent” to facilitate collaboration among researchers from different domains connector. For example, LLMs helped by translating domain-specific terminology into language that visualization researchers could understand.
@@ 4 Core: Discover
The “Discover” stage highlights the engagement with domain experts to characterize the domain problems by iteratively refining understanding through discussions. Novice researchers often used LLMs to explore the literature review by asking LLMs to suggest papers or summarize the latest research trends assistant connector, which was particularly helpful for those unfamiliar with the domain. However, this strategy often introduced hallucinated references to novices, where LLMs usually provided nonexistent papers or incorrect paper summarization. According to P21, who is a professor and lecturer in visualization, LLMs often fail to recognize the graphical visual design elements, but the students cannot totally identify this incorrectness. Proficient PhD students and professors used LLMs to summarize the takeaways from the academic paper assistant, e.g., key contributions, methodologies, and takeaways from visualization papers. However, participants commented that LLMs often misinterpret the contributions of visualization papers, particularly related to understanding visual elements. For instance, the models often failed to accurately understand visual marks and channels for either visualization idioms or novel designs. This limitation required significant additional manual effort to validate outputs.
Beyond literature discovery, participants also explored interesting ways to use LLMs to understand domain-specific requirements connector. Instead of relying solely on interviews with domain experts, some researchers used LLMs to analyze routine tasks performed by professionals in their fields. For example, P12 described using Copilot embedded in an IDE to document how bioinformatics researchers analyze tumor tissue slices on their laptops, which capture the practical workflows that might otherwise be overlooked in traditional interviews. This approach can significantly reduce information loss and provide a more precise understanding of domain requirements.
Participants also expressed that they often use LLMs to double-check the collected design requirements from domain experts connector. However, LLMs’ answers are usually too generic to help with the requirement validation. This limitation was particularly evident when researchers leveraged LLMs to generate design requirements or propose novel solutions. As a result, researchers emphasized the importance of combining LLM outputs with their own expertise, using the models as tools to augment, rather than replace, their discovery process. P14 also mentioned the usage of the generation of the questionnaire for the pilot study or formative study assistant. Also, P14 used LLMs to generate the possible answers from the provided questionnaire simulator, which is used to compare with the real answers from the participants to identify some unique and insightful findings.
@@ 5 Core: Design
The “Design” stage focuses on creating data abstractions, visual encodings, and interaction mechanisms based on the shared understanding developed in the discovery phase. Three professors highlighted the use of LLMs as “translators” during the design process to bridge the gap between visualization designers and domain experts connector. In this case, LLMs were used to facilitate communication by summarizing iterative meetings, generating meeting notes, and translating domain-specific knowledge and terminology into practical insights for visualization designers. Some PhD students reported using LLMs to decompose domain requirements, obtained during the “Discover” stage, into sub-tasks and brainstorm design ideas based on these sub-tasks assistant. For example, they would ask LLMs to search approaches used in prior visualization studies, such as those published in IEEE TVCG papers, for similar design requirements.
@@
For system-wide design tasks, such as determining the overall layout, interaction logic, and workflow between views, LLMs were used to provide high-level suggestions based on the screenshot together with the code simulator. However, participants noted that LLMs often struggled with these tasks because LLMs lacked the ability to maintain a comprehensive understanding of the global system. This limitation required researchers to manually refine and integrate the LLM’s outputs into a cohesive global design.
LLMs were also employed as sources of design inspiration assistant. Some researchers provided a domain and asked the LLM to suggest straightforward visualization designs to inspire their work, while participants noted that the suggestions were often limited to those commonly-used approaches. To address this challenge, researchers came up with practices such as incorporating evaluation metrics (e.g., significance or novelty) to assess the quality of the proposed designs. Moreover, PhD designers used LLMs to search commercial tools online for existing solutions as inspiration for their designs assistant.
@@ 6 Core: Implement
The “Implement” stage in the design study methodology explores rapidly prototyping visualization tools while remaining flexible to changes based on ongoing feedback. According to our interview, this stage is the common scenario where visualization developers rely on LLMs. First of all, for data processing tasks, novices and PhD students can use LLMs to generate code for data cleaning or other pre-processing tasks programmer. Also, after the system prototype becomes mature, LLMs can be employed to generate large-scale datasets to test the system’s scalability programmer, which can support researchers to initially evaluate if the prototype can seamlessly fit the domain’s real needs.
In terms of technical stack setup, according to the findings from the professors, LLMs can help beginners to generate basic but functional project architectures programmer, such as a simple React.js and Flask framework that supports fundamental tasks like data requesting. This approach allows developers to focus on design implementation while avoiding the time-consuming process of learning code frameworks.
For implementing the design, LLMs can assist in generating code for visual design implementation programmer. While LLMs work well for basic chart designs when provided with a sketch or bitmap, they often fail to recognize hierarchical structures. For example, in the case of a stacked bar chart, LLMs may generate separate bars instead of a cohesive stacked structure. As a workaround, novices may turn to user-friendly tools like Tableau, but it introduced new issues of flexibility and interactivity, limiting the visualization to only simple analysis tasks. Another significant challenge arises when the generated JavaScript code is only suitable for static HTML pages but fails to integrate into component-based frameworks like React.js. This incompatibility comes from several issues: 1) complex interaction logic makes the code difficult to adapt; 2) conflicts such as class name duplication can cause system bugs; 3) the generated design may not align with the overall size or color scheme of the system; and 4) SVG elements may not be correctly grouped. The best practices to address these challenges include providing detailed context about all components to hint to LLMs.
Several challenges also exist in the process of fine-tuning prototype systems. For example, it is difficult for LLMs to modify specific annotated parts of a design draft, as they struggle to recognize which part of the design is being referenced. Also, if a user requests modifying the “top left” circle, LLMs lack spatial awareness and cannot identify the target. Moreover, some descriptions, such as making a color scheme “look fancy”, are often misunderstood by LLMs. A possible practice to address these issues is to write detailed comments in the code that explicitly specify the modifications required.
@@ 7 Core: Deploy
The “Deploy” stage basically means releasing the visualization tool in a real-world context and gathering feedback from domain experts to assess its effectiveness. Advanced students have reported using LLMs to simulate user behavior, allowing the model to “pretend” to be a user simulator. This approach helps uncover potential use cases for the system, as well as identify alternative workflows that differ from conventional ones proposed by human experts. However, two significant challenges arise: 1) the generic background knowledge of LLMs can lead to superficial stories that lack depth;and 2) hallucinations by LLMs may result in incorrect stories based on non-existent system features.
@@
Moreover, proficient students have mentioned some useful ways to support user studies and feedback analysis by LLMs. Specifically, developers can transcribe user study recordings into text and use LLMs to summarize key takeaways and feedback assistant. Similarly, LLMs can assist in generating questionnaires and designing tasks based on the provided system workflow assistant. However, challenges exist in these practices. For example, relying on LLMs to summarize feedback may result in conclusions that do not align with the developer’s domain expertise. Additionally, when generating questionnaires based on prior papers, LLMs may over-rely on the provided materials, leading to repetitive questions.